Suggested time: 60 minutes. Answer both questions. 1. Consider the following balance sheet for DeMontfort Savings Bank plc. (in millions): Liabilities and Equity Demand deposits Assets Floating-rate mortgages (currently 10% annually) 30-year fixed-rate loans (currently 7% annually) E70 £50__ (currently 6%,annually). Time deposits (currently 6% annually) £20 £10 Total Liabilities & Equity £100 £50 Equity Total Assets £100 What is DeMontfort Savings Banks expected net interest income (a) at year-end? (3 marks) (b) What will be the net interest income at year-end if interest rates rise by 2 percent? (3 marks) (c) Using the cumulative repricing gap model, what is the expected net interest income for a 2 percent increase in interest rates? (3 marks) (d) What will be the net interest income at year-end if interest rates increase 200 basis points on assets, but only 100 basis points on liabilities? Is it reasonable for changes in interest rates to affect balance sheet in an uneven manner? Why? (6 marks) PLEASE TURN OVER
Answer #1

(a)

Current expected interest income: 50m(0.10) + 50m(0.07) = 8.5m.

Expected interest expense:             70m(0.06) ‑ 20m(0.07) = 5.6m.

Expected net interest income:                          8.5m ‑ 5.6m = 2.9m

(b)

After the 200 basis point interest rate increase (or 2%), net interest income declines to:

         50(0.12) + 50(0.07) ‑ 70(0.08) ‑ 20(.07) = 9.5m ‑ 7.0m = 2.5m, thus a decline of 0.4m. (2.9m - 2.5m)

(c)

DeMontfort's repricing or funding gap is 50m ‑ 70m = ‑20m. The change in net interest income using the cumulative funding gap model is (‑20m)(0.02) = ‑$.4m.

(d)

After the unequal rate increases, net interest income will be 50(0.12) + 50(0.07) ‑ 70(0.07) ‑ 20(0.07) = 9.5m ‑ 6.3m = 3.2m, an increase of 0.3m. It is not uncommon for interest rates to adjust in an unequal manner on assets versus liabilities. Interest rates often do not adjust solely because of market pressures. In many cases the changes are affected by decisions of management and other internal overriding factors.